Circular No. 8940

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 J1 (BOATTINI) A. Boattini reports his discovery of a comet with a coma diameter of

∼ 50′′ and a fan-shaped tail 2′ long in p.a. 235^{o} on Catalina Sky Survey CCD images obtained with the 0.68-m Schmidt telescope (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, several other CCD astrometrists have found this object to show cometary appearance. S. Nakano, Sumoto, Japan, reports that images obtained on May 2.73–2.74 UT by A. Asami and T. Sakamoto (Bisei Spaceguard Center 1.0-m f/3 reflector) show a coma diameter of 20′′ and a tail 24′′ long in p.a. 230^{o}. R. Apitzsch (Wildberg, Germany, 0.35-m f/4.2 reflector, May 3.02–3.03) finds a 1′ tail toward the southwest. V. Gonano, E. Guido, and G. Sostero (Remanzacco, Italy, 0.45-m f/4.4 reflector) report that thirty stacked 60-s unfiltered exposures obtained during May 3.02–3.05 show a compact coma ∼ 15′′ in diameter (total mag ∼ 14.7) with a central condensation and an apparent tail 35′′ long in p.a. 280^{o}. R. Ligustri (Talmassons, Udine, Italy, 0.35-m f/5 reflector, May 3.08–3.09) writes that thirty 60-s exposures show a faint tail 45′′ long. J. M. Aymami (Tiana, Spain, 0.25-m f/5 reflector, May 3.11–3.12) finds a nuclear condensation and a short, fan-like tail ≈ 16′′ long in p.a. 238^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
May 2.46097  21^{h}01^{m}39.^{s}57  +10^{o}19′29′′.9  14.2

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-J17.

T = 2008 Mar. 17.336 TT  ω = 0.^{o}915
Ω = 257.241 }2000.0
q = 1.54922 AU  i = 46.374

V5579 SAGITTARII = NOVA SAGITTARII 2008 N. N. Samus, Institute of Astronomy, Moscow, informs us that the

nova announced on IAUC 8937 is being assigned the variable-star designation V5579 Sgr. Additional photometry from M. M. M. Santangelo, Capannori, Italy (cf. CBET 1352): Apr. 27.124 UT, V = 7.56 ± 0.04, B–V = +1.33 ± 0.05. Visual magnitude estimates: Apr. 24.083, 6.7 (K. Hornoch, Vohancice, Czech Republic); 25.35, 6.9 (W. Souza, Sao Paulo, Brazil; moonlight); 26.29, 7.8 (R. Fidrich, Bakonycsernye, Hungary); 27.361, 7.7 (Souza); 28.361, 8.0 (Souza).

2008 May 3  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green